OpenStudy (anonymous):

HELP PLEASE!! I Will Give Metals.. Write an equation for the circle with center (-4, 4) and containing the point (-1, 16).

OpenStudy (kohai):

(x-h)^2 + (y-k)^2 = r^2 (h,k) being the center of the circle and r being the radius you're given (-4,4) as your (h,k) so plugging that into the equation you get (x+4)^2 + (y-4)^2 = r^2 you are still missing your "r" (-1,16) is (x,y) so plugging that in for for x and y you'll have (-1+4)^2 + (16-4)^2 = r^2 (3)^2 + (12)^2 = r^2 9 + 144 = r^2 153 = r^2 sqrt 153 = r^2 and plug it all in (x+4)^2 + (y-4)^2 = sqrt 153 and you're all done :)

OpenStudy (anonymous):

Thank you so much!

OpenStudy (kohai):

You're welcome :)